# VoxEval: Benchmarking the Knowledge Understanding Capabilities of End-to-End Spoken Language Models

### v2

@@ 4.2 Results
We aim to answer the following Research Questions (RQs) through experiments. RQ1: What is the overall performance of existing SLMs on the VoxEval benchmark? RQ2: How do existing SLMs perform across various knowledge domains? RQ3: To what extent do different input audio conditions influence the performance of existing SLMs? RQ4: How well are current SLMs in reasoning, particularly in mathematical reasoning?
@@
To answer RQ1, we evaluate the five SLMs on VoxEval, whose results are shown in Table 3. We observe that existing SLMs perform poorly on VoxEval. In fact, most SLMs’ performance does not even surpass random guessing. Since each question in VoxEval offers four answer choices, the expected score for random guessing is 0.25. However, all SLMs, except for GLM-4-Voice, fail to exceed this baseline. This indicates that, in many instances, the SLMs struggle to correctly follow instructions to select from the options. This highlights the difficulty of enabling SLMs to “speak" the correct answer, making VoxEval a particularly challenging benchmark for current SLMs. A case study of different SLMs’ spoken responses is shown in Appendix B. During the evaluation, we found that SpeechGPT’s performance is close to 0%. This is likely due to our approach of directly performing speech-to-speech generation without utilizing the "chain-of-modality" mode described in Zhang et al. (2023). We choose not to use the "chain-of-modality" approach because it involves converting the input into text before generating a response, which we believe would significantly increase inference latency and is not an optimal strategy for end-to-end SLMs.
@@
To answer RQ2, we present the SLM performance results across various subject groups within VoxEval. Following MMLU, all knowledge subjects are divided into four groups: STEM (18 subjects), Social Science (12 subjects), Humanities (13 subjects), and Others (13 subjects). Table 4 shows the average performance of the subjects in each group. We notice that SLMs exhibit significant performance variations across different subject groups. For instance, when using alloy as the input voice, SPIRIT-LM shows a score difference of approximately 4%, and GLM-4-Voice demonstrates a difference of around 10%. Moreover, different SLMs excel in different subject groups. For example, TWIST, Moshi, and GLM-4-Voice typically perform the best in the Others group, STEM group, and Social Science group, respectively.
@@
To answer RQ3, we present the overall performance of SLMs across various input audio conditions in Table 3. Our key findings are summarized as follows.
@@
  • SLMs are susceptible to different input audio conditions. Although the performance of SLMs does not vary significantly with changes in speaker voices, we observe that differences in speaking styles and audio quality can negatively impact their performance, sometimes leading to a performance drop of up to 6%.
@@
We further visualize performance differences across all the settings (input audio conditions), as depicted in Figure 2. For each setting, we compute the maximum performance difference of all 56 subjects in the VoxEval dataset and present the results (56 difference scores) using box plots. In the “Speaker" setting, the differences are calculated using audio from all six speakers provided by OpenAI TTS. For the remaining settings, since all audio variations are applied to speech data generated by the “alloy" voice, the differences are computed by comparing the original “alloy" audio with the transformed audio. The top part of the figure illustrates the absolute differences in scores, while the bottom part shows the relative percentage differences. We have the following key observations.
@@
  • Models exhibit varying levels of stability when faced with changes in audio conditions. As shown in Figure 2(b), TWIST and Moshi tend to be less stable, whereas GLM-4-Voice demonstrates minimal relative score differences.
Additionally, SLMs display relatively consistent behavior across different condition changes. According to Table 4, when input audio conditions are altered, TWIST, Moshi, and GLM-4-Voice consistently perform the best in others subjects, STEM subjects, and social science subjects, respectively.
To answer RQ4, we present the performance of SLMs across three math subjects in VoxEval, as shown in Table 5. The selected subjects—elementary math, high school math, and college math—are intended to represent varying levels of reasoning complexity. We observe that existing SLMs perform similarly across different levels of math questions. Interestingly, in some cases, SLMs perform better on more challenging questions. For example, TWIST and GLM-4-Voice achieve the highest performance in college math. This counterintuitive fact suggests that existing SLMs are unlikely to possess basic reasoning abilities.
@@ 5 Conclusion
In this paper, we introduce VoxEval, a novel speech-based question-answering benchmark designed to evaluate the knowledge understanding capabilities of end-to-end Spoken Language Models (SLMs) under diverse audio conditions. Comprehensive evaluations of recent SLMs reveal significant performance limitations, including poor accuracies that are often below random guessing, heightened susceptibility to variations in input audio conditions, and notable challenges in domains such as mathematical reasoning. These findings highlight the need for future research to improve the robustness and reasoning abilities of SLMs in real-world conversational settings. VoxEval serves as a critical step toward advancing the evaluation and development of SLMs, paving the way for more effective and reliable speech-based AI systems.
@@ Appendix A Explanation for Various Speech Recording Conditions
Table 6 provides the input audio conditions categorized under the "other environment acoustics" and the transformation settings in the audioaugmentation package.
@@
Bitcrush Bitcrush reduces the resolution (bit depth) and/or sampling rate of a sound, creating a lo-fi, distorted, or "crunchy" effect by introducing digital artifacts and aliasing. min_bit_depth=5, max_bit_depth=6
@@
Clipping Distortion Clipping distortion occurs when an audio signal exceeds the maximum amplitude that a system can handle, causing the waveform to be "clipped" at the peaks and resulting in a harsh, distorted sound. min_percentile_threshold=20, max_percentile_threshold=20
@@
Table 6: Explanations and transformation settings categorized under "other environment acoustics."
Appendix B Case Study for SLM Responses
@@
GLM-4-Voice The correct answer is B. Multiplication is not associative. GLM-4-Voice is able to answer with required format as well as giving an explanation afterwards.

### v3

@@ 4.2 Results
We aim to answer the following Research Questions (RQs) through experiments. RQ1. What is the overall performance of existing SLMs on the VoxEval benchmark? RQ2. How do existing SLMs perform across various knowledge domains? RQ3. To what extent do different input audio conditions influence the performance of existing SLMs? RQ4. How well are current SLMs in reasoning, particularly in mathematical reasoning?
@@
To answer RQ1, we evaluate five SLMs on VoxEval, as shown in Table 5. Results reveal that current SLMs perform poorly, with most failing to surpass random guessing. Since VoxEval questions have four answer choices, random guessing yields an expected score of 25%. However, only GLM-4-Voice exceeds this baseline, indicating that SLMs often struggle to follow instructions and select the correct answer. This underscores the challenge of enabling SLMs to “speak" accurate responses, making VoxEval a tough benchmark. A case study of SLMs’ spoken responses is in Appendix C. Notably, SpeechGPT’s performance is near 0%. This is likely due to our approach of directly performing speech-to-speech generation without utilizing the “chain-of-modality" mode described in Zhang et al. (2023). We choose not to use the “chain-of-modality" approach because it involves converting the input into text before generating a response, which we believe would significantly increase inference latency and is not an optimal strategy for end-to-end SLMs.
Since some SLMs support both speech and text as input/output (I/O), we further evaluate the performance of SLM when using the four different I/O modality combinations. In particular, we focus on SPIRIT-LM, as it is the only model that supports both modalities for input and output during inference. The evaluation results, shown in Table 6, reveal that SPIRIT-LM’s performance varies significantly depending on the modality combination used, with differences exceeding 20%. This highlights the difficulty of achieving consistent behavior across various I/O configurations.
@@
To answer RQ2, we present the SLM performance results across various subject groups within VoxEval. Following MMLU, all knowledge subjects are divided into four groups: STEM (18 subjects), social science (12 subjects), humanities (13 subjects), and others (13 subjects). Table 7 shows the average performance of the subjects in each group. We notice that SLMs exhibit significant performance variations across different subject groups. For instance, when using alloy as the input voice, SPIRIT-LM shows a score difference of approximately 4%, and GLM-4-Voice demonstrates a difference of around 10%. Moreover, different SLMs excel in different subject groups. For example, despite the change in input audio conditions in various settings, TWIST, Moshi, and GLM-4-Voice typically perform the best in the others, STEM, and social science groups, respectively.
@@
To answer RQ3, we present the overall performance of SLMs across various input audio conditions in Table 5. Our key findings include:
@@
  • SLMs are susceptible to different input audio conditions. Although SLM performances do not vary significantly across different speaker voices, variations in speaking styles and audio quality can cause up to a 6% performance drop.
@@
We further visualize performance differences across all input audio condition settings, as shown in Figure 2. For each setting, we compute the maximum score difference for each of the 56 subjects in the VoxEval dataset, based on the scores obtained across all results in the corresponding setting. This yields 56 difference scores for each setting, which are presented using box plots. In the “Speaker" setting, the differences are calculated using audio from all six speakers provided by OpenAI TTS. For the remaining settings, since all audio variations are applied to speech data generated by the “alloy" voice, the differences are computed by comparing the scores of the original “alloy" audio with those of the transformed audio. The top part of the figure illustrates the absolute differences in scores, while the bottom part shows the relative percentage differences. Algorithm 1 presents the pseudo-code for generating the box plots. We make the following key observations.
@@
  • Models exhibit varying levels of stability when faced with changes in audio conditions. Lower score differences indicate better stability. As shown in Figure 2(b), TWIST and Moshi tend to be less stable, whereas GLM-4-Voice demonstrates minimal relative score differences.
To answer RQ4, we present the performance of SLMs across three math subjects in VoxEval, as shown in Table 8. The selected subjects—elementary math, high school math, and college math—are intended to represent varying levels of reasoning complexity. The math-solving abilities of SLMs are tested under two scenarios: direct answer, where the model outputs the answer directly, and Chain-of-Thought prompting (CoT) Wei et al. (2022), which uses annotated intermediate steps from three dev-set questions per subject as examples appended before the actual question. All the CoT examples and questions are in speech format. CoT evaluations are conducted in two settings: Cut, where the input audio is trimmed to the last 80 seconds (consistent with other experiments), and No cut, where the audio is left untrimmed to preserve longer CoT prompts. Two metrics are used for evaluation: string matching, the standard final answer extraction method, and LLM, where GPT-4o Hurst et al. (2024) is used to extract answers to complement string matching, as models often include lengthy reasoning steps that hinder accurate string-based extraction when applying CoT. Our observations include:
  • Existing SLMs perform similarly across different levels of math questions. Interestingly, in some cases, SLMs perform better on more challenging questions. For example, TWIST and GLM-4-Voice achieve the highest performance in college math. This counterintuitive fact suggests that existing SLMs are unlikely to possess basic reasoning abilities.
  • Using CoT reduces the performance of SLMs. We observe this trend consistently across all tested models. This suggests that CoT does not enhance SLMs in the same way it improves TLMs, emphasizing the need for specialized techniques to improve SLM reasoning.
  • Extending CoT example length does not significantly impact SLM performances. Most models perform similarly in Cut and No cut scenarios, but TWIST struggles in No cut, suggesting some models are sensitive to input audio length.
@@ 5 Conclusion
This paper introduces VoxEval, a speech question-answering benchmark for assessing the knowledge understanding of end-to-end Spoken Language Models (SLMs) across diverse audio conditions. Evaluations of recent SLMs reveal significant limitations, including low accuracies that are often below random guessing, sensitivity to variations in input audio conditions, and difficulties with reasoning tasks like mathematical problem-solving. These results emphasize the need for further research to improve SLMs’ knowledge understanding, robustness, and reasoning capabilities in real-world conversational settings. VoxEval serves as a critical step toward advancing the evaluation and development of SLMs, paving the way for more effective and reliable speech-based AI systems.
6 Limitations
The limitations of this work come from the scope and the evaluations. First, the VoxEval benchmark focuses exclusively on assessing the robustness of SLMs in terms of knowledge understanding, leaving out other critical aspects such as fairness, toxicity, hallucination, etc. These issues are particularly relevant in generative scenarios, where SLMs may exhibit biases, generate harmful content, or produce inaccurate information. However, these aspects fall outside the scope of VoxEval and require separate evaluation frameworks tailored to address them comprehensively. Second, the evaluation of the four modality combinations (speech-to-speech, speech-to-text, text-to-speech, and text-to-text) is not comprehensive enough, as it only includes one SLM (SPIRIT-LM). Expanding this evaluation to a broader range of models would enhance the reliability and generalizability of the findings.
@@ Appendix A Explanation for Various Speech Recording Conditions
Table 9 provides the input audio conditions categorized under the “other environment acoustics" and the transformation settings in the audioaugmentation package.
@@
Bitcrush Bitcrush reduces the resolution (bit depth) and/or sampling rate of a sound, creating a lo-fi, distorted, or “crunchy" effect by introducing digital artifacts and aliasing. min_bit_depth=5, max_bit_depth=6
@@
Clipping Distortion Clipping distortion occurs when an audio signal exceeds the maximum amplitude that a system can handle, causing the waveform to be “clipped" at the peaks and resulting in a harsh, distorted sound. min_percentile_threshold=20, max_percentile_threshold=20
@@
Table 9: Explanations and transformation settings categorized under “other environment acoustics."
Appendix B Additional Experimental Settings
For models used in our evaluation, SpeechGPT, TWIST, SPIRIT-LM, and Moshi contain 7B parameters, and GLM-4-Voice contains 9B parameters. We use the base version of SPIRIT-LM. All experiments are conducted on a single NVIDIA RTX A6000 GPU. For all the models, we use the default hyperparameters provided in their code implementations. Each experiment is performed once.
Appendix C Case Study for SLM Responses
@@
GLM-4-Voice The correct answer is B. Multiplication is not associative. GLM-4-Voice can respond in the specified format and provide an explanation afterward, though the explanation may occasionally be inaccurate.
@@
Appendix D String Matching vs. LLM Extraction
In Table 8, we notice that in most cases, the accuracy of using an LLM to extract the final answer is lower. This is primarily due to edge cases. For instance, when the SLM generates an output like
"is actually to where h or 0 0 1 times 2 squared. Option D, to where v is a positive, moving by 1 times 2 squared. Statement 1, because of the path grow entertained by t is",
the string matching algorithm identifies D as the final answer. However, the SLM does not explicitly state that D is the correct answer. In such situations, the LLM produces “None!" as the output. If D happens to be the correct answer, string matching would classify it as correct, whereas the LLM would mark it as incorrect. The full prompt used for extracting the final answer with the LLM is provided in Figure 5.
@@
{SLM output}
# Instruction
Above is the answer provided by an AI model for a Multiple Choice Question with four answer choices (A, B, C, or D). Based on the above text, extract the final answer from it. If you can find the answer from the above text, only output the answer choice. Do not include anything else in your output. For example, a possible output is "C". If you cannot find the answer, only output "None!".
Figure 5: The prompt for GPT-4o to extract the final answers from SLM outputs.
Algorithm 1 Procedure for Generating Box Plots of Performance Differences
1:  Input: VoxEval dataset with $N=56$ subjects, input audio condition settings $\{S_{1},S_{2},\dots,S_{k}\}$, and corresponding groups of results for each setting.
2:  Output: Box plots visualizing performance differences for each setting.
3:  for each input audio condition setting $S_{i}$ do
4:     Extract all result groups $\{G_{1},G_{2},\dots,G_{m}\}$ corresponding to $S_{i}$.
5:     Initialize an empty list $D_{i}=[]$ to store difference scores for $S_{i}$.
6:     for each subject $n\in\{1,2,\dots,N\}$ do
7:        Collect scores $\{s_{n,1},s_{n,2},\dots,s_{n,m}\}$ for subject $n$ across all $m$ groups in $S_{i}$.
8:        Compute the maximum absolute score difference:
$d_{n}=\max_{j,k}\big{|}s_{n,j}-s_{n,k}\big{|}\quad\text{for }j,k\in\{1,2,\dots,m\}.$
9:        Append $d_{n}$ to $D_{i}$: $D_{i}\leftarrow D_{i}\cup\{d_{n}\}$.
10:     end for
11:     Use $D_{i}=\{d_{1},d_{2},\dots,d_{56}\}$ to generate a box plot for $S_{i}$.
12:  end for
13:  Output: Box plots for all input audio condition settings $\{S_{1},S_{2},\dots,S_{k}\}$.